# Step-by-Step Diffusion: An Elementary Tutorial

### v1

@@ Acknowledgements
We are grateful for helpful feedback and suggestions from many people, in particular: Josh Susskind, Eugene Ndiaye, Dan Busbridge, Sam Power, De Wang, Russ Webb, Sitan Chen, Vimal Thilak, Etai Littwin, Chenyang Yuan, Alex Schwing, and Miguel Angel Bautista Martin.
@@ 1.1 Gaussian Diffusion
The key insight of diffusion is, learning to reverse each intermediate step can be easier than learning to sample from the target distribution in one step555 Intuitively this is because the distributions $(p_{t-1},p_{t})$ are already quite close, so the reverse sampler does not need to do much.. There are many ways to construct reverse samplers, but for concreteness let us first see the standard diffusion sampler which we will call the DDPM sampler [^6] This is the sampling strategy originally proposed in Sohl-Dickstein et al. (2015)..k
@@ 2.4 Diffusions as SDEs [Optional]
In this section, we connect the discrete-time processes we have discussed so far to stochastic differential equations (SDEs). In the continuous limit, as ${\Delta{t}}\to 0$, our discrete diffusion process turns into a stochastic differential equation. SDEs can also represent many other diffusion variants (corresponding to different drift and diffusion terms), offering flexibility in design choices, like scaling and noise-scheduling. The SDE perspective is powerful because existing theory provides a general closed-form solution for the time-reversed SDE. Discretization of the reverse-time SDE for our particular diffusion immediately yields the sampler we derived in this section, but reverse-time SDEs for other diffusion variants are also available automatically (and can then be solved with any off-the-shelf or custom SDE solver), enabling better training and sampling strategies as we will discuss further in Section 5. Though we mention these connections only briefly here, the SDE perspective has had significant impact on the field. For a more detailed discussion, we recommend Yang Song’s blog post (Song, 2021).
@@
$x_{t-{\Delta{t}}}$ $=x_{t}-\frac{{\Delta{t}}}{t}\mathop{\mathbb{E}}[x_{0}-x_{t}\mid x_{t}]+\mathcal{N}(0,\sigma_{q}^{2}{\Delta{t}})$
[…]
@@ 3 Deterministic Sampling: DDIM
then we want to show that232323 The notation $F\sharp p$ means the distribution of $\{F(x)\}_{x\sim p}$. This is called the pushforward of $p$ by the function $F$.
@@ 3.1 Case 1: Single Point
The correctness of Algorithm 2 still holds272727 See Claim 5 in Appendix B.3 for an explicit statement. if $x_{0}$ is an arbitrary point instead of $x_{0}=0$, since everything is transitionally symmetric.
@@ 3.5 Discussion: DDPM vs DDIM
DDPM will ideally produce a sample from $p(x_{0}\mid x_{1})$. If the forward process mixes sufficiently (i.e. for large $\sigma_{q}$ in our setup), then the final point $x_{1}$ will be nearly independent from the initial point. Thus $p(x_{0}\mid x_{1})\approx p(x_{0})$, so the distribution output by the ideal DDPM will not depend at all363636 Actual DDPMs may have a small dependency on the initial point $x_{1}$, because they do not mix perfectly (i.e. the final distribution $p_{1}$ is not perfectly Gaussian). Randomizing the initial point may thus help with sample diversity in practice. on the starting point $x_{1}$. In contrast, DDIM is deterministic, so it will always produce a fixed value for a given $x_{1}$, and thus will depend very strongly on $x_{1}$.
@@ 3.6 Remarks on Generalization
In this tutorial, we have not discussed the learning-theoretic aspects of diffusion models: How do we learn properties of the underlying distribution, given only finite samples and bounded compute? These are fundamental aspects of learning, but are not yet fully understood for diffusion models; it is an active area of research373737 We recommend the introductions of Chen et al. (2022) and Chen et al. (2024b) for an overview of recent learning-theoretic results. This line of work includes e.g. De Bortoli et al. (2021); De Bortoli (2022); Lee et al. (2023); Chen et al. (2023, 2024a). .
To appreciate the subtlety here, suppose we learn a diffusion model using the classic strategy of Empirical Risk Minimization (ERM): we sample a finite train set from the underlying distribution, and optimize all regression functions w.r.t. this empirical distribution. The problem is, we should not perfectly minimize the empirical risk, because this would yield a diffusion model which only reproduces the train samples383838 This is not specific to diffusion models: any perfect generative model of the empirical distribution will always output a uniformly random train point, which is far-from-optimal w.r.t. the true underlying distribution..
@@ 4 Flow Matching
We now introduce the framework of flow matching (Lipman et al., 2023; Albergo et al., 2023; Liu et al., 2022). Flow matching can be thought of as a generalization of DDIM, which allows for more flexibility in designing generative models— including for example the rectified flows (sometimes called linear flows) used by Stable Diffusion 3 (Liu et al., 2022; Esser et al., 2024).
@@ 4.1 Flows
Just as flows define maps between initial and final points, they also define transports between entire distributions, by “pushing forward” points from the source distribution along their trajectories. If $p_{1}$ is a distribution on initial points [^39] Notational warning: Most of the flow matching literature uses a reversed time convention, so $t=1$ is the target distribution. We let $t=0$ be the target distribution to be consistent with the DDPM convention., then applying the flow $v$ yields the distribution on final points404040 We could equivalently write this as the pushforward $\mathrm{RunFlow}(v,\cdot,0)\sharp p_{1}$.
@@ 4.3 Marginal Flows
We would like to combine all of these pointwise flows somehow, to get a single flow $v^{*}$ that implements the same transport between distributions444444 Why would we like this? As we will see later, it simplifies our learning problem: instead of having to learn the distribution of all the individual trajectories, we can instead just learn one velocity field representing their bulk evolution.. Our previous discussion [^45] Compare to Equation (47) in Section 3. A formal statement of how to combine flows is given in Appendix B.4. in Section 3 tells us how to do this: to determine the effective velocity $v_{t}^{*}(x_{t})$, we should take a weighted-average of all individual particle velocities $v^{[x_{1},x_{0}]}_{t}$, weighted by the probability that a particle at $x_{t}$ was generated by the pointwise flow $v^{[x_{1},x_{0}]}$. The final result is464646 An alternate way of viewing this result at a high level is: we start with pointwise flows $v^{[x_{1},x_{0}]}$ which transport delta distributions: $\delta_{x_{1}}\overset{v^{[x_{1},x_{0}]}}{\hookrightarrow}\delta_{x_{0}}.$ (60) And then Equation (62) gives us a fancy way of “averaging these flows over $x_{1}$ and $x_{0}$”, to get a flow $v^{*}$ transporting $q=\mathop{\mathbb{E}}_{x_{1}\sim q}[\delta_{x_{1}}]\overset{v^{*}}{\hookrightarrow}\mathop{\mathbb{E}}_{x_{0}\sim p}[\delta_{x_{0}}]=p.$ (61)
@@ 4.4 A Simple Choice of Pointwise Flow
We need an explicit choices of: pointwise flow, base distribution $q$, and coupling $\Pi_{q,p}$. There are many simple choices which would work474747 Diffusion provides one possible construction, as we will see later in Section 4.6..
@@
which simply linearly interpolates between $x_{1}$ and $x_{0}$ (and corresponds to the choice made in Liu et al. (2022)). In Figure 5 we visualize a marginal flow composed of linear pointwise flows, the same annular base distribution $q$ of Figure 4, and target distribution equal to a point-mass ($p=\delta_{x_{0}}$) [^48] A marginal distribution with a point-mass target distribution – or equivalently the average of pointwise flows over the the base distribution only – is sometimes called a (one-sided) conditional flow (Lipman et al., 2023). .
@@ 4.5 Flow Matching
    Compute the point $x_{t}$ deterministically, by running505050 If we chose linear pointwise flows, for example, this would mean $x_{t}\leftarrow tx_{1}+(1-t)x_{0}$, via Equation (64). the pointwise flow $v^{[x_{1},x_{0}]}$ starting from point $x_{1}$ up to time $t$.
@@
To sample from the trained model (our estimate of $v_{t}^{*}$), we first sample a source point $x_{1}\sim q$, then transport it along the learnt flow to a target sample $x_{0}$. Pseudocode listings  4 and  5 give the explicit procedures for training and sampling from flow-based models (including the special case of linear flows for concreteness; matching Algorithm 1 in Liu et al. (2022).).
@@ 4.6 DDIM as Flow Matching [Optional]
    which both generate the trajectory515151 See Appendix B.6 for details on why  (68) and (69) are equivalent along their trajectories.:
@@
This claim is straightforward to prove (see Appendix B.5), but the implication is somewhat surprising: we can recover the DDIM trajectories (which are not straight in general) as a combination of the straight pointwise trajectories in Equation (70). In fact, the DDIM trajectories are exactly equivalent to flow-matching trajectories for the above linear flows, with a different scaling of time ($\sqrt{t}$ vs. $t$)525252 DDIM at time $t$ corresponds to the linear flow at time $\sqrt{t}$; thus linear flows are “slower” than DDIM when $t$ is small. This may be beneficial for linear flows in practice (speculatively).. {mdframed}[nobreak]
@@ 4.7 Additional Remarks and References [Optional]
  • What we called pointwise flows are also called two-sided conditional flows in the literature, and was developed in Albergo and Vanden-Eijnden (2022); Pooladian et al. (2023); Liu et al. (2022); Tong et al. (2023).
@@ 5 Diffusion in Practice
One advantage of the VAE interpretation is, it gives us an estimate of the data likelihood under our generative model, by using the standard Evidence-Based-Lower-Bound (ELBO) for VAEs. This allows us to train diffusion models directly using a maximum-likelihood objective. It turns out that the ELBO for the diffusion VAE reduces to exactly the L2 regression loss that we presented, but with a particular time-weighting that weights the regression loss differently at different time-steps $t$. For example, regression errors at large times $t$ (i.e. at high noise levels) may need to be weighted differently from errors at small times, in order for the overall loss to properly reflect a likelihood.575757 See also Equation (5) in Kadkhodaie et al. (2024) for a simple bound on KL divergence between the true distribution and generated distribution, in terms of regression excess risks. The best choice of time-weighting in practice, however, is still up for debate: the “principled” choice informed by the VAE interpretation does not always produce the best generated samples585858 For example, Ho et al. (2020) drops the time-weighting terms, and just uniformly weights all timesteps.. See Kingma and Gao (2023) for a good discussion of different weightings and their effect.
@@ B.1 KL Error in Gaussian Approximation of Reverse Process
$=F(0)+0.5\sigma^{2}(F^{\prime\prime}(0)-F^{\prime}(0)^{2})+O(\sigma^{4})$ (by Line 101)
@@ B.6 Two Pointwise Flows for DDIM give the same Trajectory
We want to show that pointwise flow 69:
@@ B.8 Proof Sketch of Claim 2
Now continuing from Line 127,

### v2

@@ Acknowledgements
We are grateful for helpful feedback and suggestions from many people, in particular: Josh Susskind, Eugene Ndiaye, Dan Busbridge, Sam Power, De Wang, Russ Webb, Sitan Chen, Vimal Thilak, Etai Littwin, Chenyang Yuan, Alex Schwing, Miguel Angel Bautista Martin, and Dilip Krishnan.
@@ 1.1 Gaussian Diffusion
The key insight of diffusion is, learning to reverse each intermediate step can be easier than learning to sample from the target distribution in one step555 Intuitively this is because the distributions $(p_{t-1},p_{t})$ are already quite close, so the reverse sampler does not need to do much.. There are many ways to construct reverse samplers, but for concreteness let us first see the standard diffusion sampler which we will call the DDPM sampler [^6] This is the sampling strategy originally proposed in Sohl-Dickstein et al. (2015)..
@@ 2.4 Diffusions as SDEs [Optional]
In this section [^19] Sections marked “[Optional]” are advanced material, and can be skipped on first read. None of the main sections depend on Optional material., we connect the discrete-time processes we have discussed so far to stochastic differential equations (SDEs). In the continuous limit, as ${\Delta{t}}\to 0$, our discrete diffusion process turns into a stochastic differential equation. SDEs can also represent many other diffusion variants (corresponding to different drift and diffusion terms), offering flexibility in design choices, like scaling and noise-scheduling. The SDE perspective is powerful because existing theory provides a general closed-form solution for the time-reversed SDE. Discretization of the reverse-time SDE for our particular diffusion immediately yields the sampler we derived in this section, but reverse-time SDEs for other diffusion variants are also available automatically (and can then be solved with any off-the-shelf or custom SDE solver), enabling better training and sampling strategies as we will discuss further in Section 5. Though we mention these connections only briefly here, the SDE perspective has had significant impact on the field. For a more detailed discussion, we recommend Yang Song’s blog post (Song, 2021).
@@
$x_{t}-x_{t-{\Delta{t}}}$ $=-\frac{{\Delta{t}}}{t}\mathop{\mathbb{E}}[x_{0}-x_{t}\mid x_{t}]+\mathcal{N}(0,\sigma_{q}^{2}{\Delta{t}})$ (31)
$=-\mathop{\mathbb{E}}[x_{t-{\Delta{t}}}-x_{t}\mid x_{t}]+\mathcal{N}(0,\sigma_{q}^{2}{\Delta{t}})$ (by Eqn. 23)
$\implies x_{t-{\Delta{t}}}$ $=\mathop{\mathbb{E}}[x_{t-{\Delta{t}}}\mid x_{t}]+\mathcal{N}(0,\sigma_{q}^{2}{\Delta{t}})$ (32)
@@ 3 Deterministic Sampling: DDIM
then we want to show that242424 The notation $F\sharp p$ means the distribution of $\{F(x)\}_{x\sim p}$. This is called the pushforward of $p$ by the function $F$.
@@ 3.1 Case 1: Single Point
The correctness of Algorithm 2 still holds282828 See Claim 5 in Appendix B.3 for an explicit statement. if $x_{0}$ is an arbitrary point instead of $x_{0}=0$, since everything is transitionally symmetric.
@@ 3.5 Discussion: DDPM vs DDIM
DDPM will ideally produce a sample from $p(x_{0}\mid x_{1})$. If the forward process mixes sufficiently (i.e. for large $\sigma_{q}$ in our setup), then the final point $x_{1}$ will be nearly independent from the initial point. Thus $p(x_{0}\mid x_{1})\approx p(x_{0})$, so the distribution output by the ideal DDPM will not depend at all373737 Actual DDPMs may have a small dependency on the initial point $x_{1}$, because they do not mix perfectly (i.e. the final distribution $p_{1}$ is not perfectly Gaussian). Randomizing the initial point may thus help with sample diversity in practice. on the starting point $x_{1}$. In contrast, DDIM is deterministic, so it will always produce a fixed value for a given $x_{1}$, and thus will depend very strongly on $x_{1}$.
@@ 3.6 Remarks on Generalization
In this tutorial, we have not discussed the learning-theoretic aspects of diffusion models: How do we learn properties of the underlying distribution, given only finite samples and bounded compute? These are fundamental aspects of learning, but are not yet fully understood for diffusion models; it is an active area of research383838 We recommend the introductions of Chen et al. (2022) and Chen et al. (2024b) for an overview of recent learning-theoretic results. This line of work includes e.g. De Bortoli et al. (2021); De Bortoli (2022); Lee et al. (2023); Chen et al. (2023, 2024a). .
To appreciate the subtlety here, suppose we learn a diffusion model using the classic strategy of Empirical Risk Minimization (ERM): we sample a finite train set from the underlying distribution, and optimize all regression functions w.r.t. this empirical distribution. The problem is, we should not perfectly minimize the empirical risk, because this would yield a diffusion model which only reproduces the train samples393939 This is not specific to diffusion models: any perfect generative model of the empirical distribution will always output a uniformly random train point, which is far-from-optimal w.r.t. the true underlying distribution..
@@ 4 Flow Matching
We now introduce the framework of flow matching (Peluchetti, 2022; Liu et al., 2022b, a; Lipman et al., 2023; Albergo et al., 2023). Flow matching can be thought of as a generalization of DDIM, which allows for more flexibility in designing generative models— including for example the rectified flows (sometimes called linear flows) used by Stable Diffusion 3 (Liu et al., 2022a; Esser et al., 2024).
@@ 4.1 Flows
Just as flows define maps between initial and final points, they also define transports between entire distributions, by “pushing forward” points from the source distribution along their trajectories. If $p_{1}$ is a distribution on initial points [^40] Notational warning: Most of the flow matching literature uses a reversed time convention, so $t=1$ is the target distribution. We let $t=0$ be the target distribution to be consistent with the DDPM convention., then applying the flow $v$ yields the distribution on final points414141 We could equivalently write this as the pushforward $\mathrm{RunFlow}(v,\cdot,0)\sharp p_{1}$.
@@ 4.3 Marginal Flows
We would like to combine all of these pointwise flows somehow, to get a single flow $v^{*}$ that implements the same transport between distributions454545 Why would we like this? As we will see later, it simplifies our learning problem: instead of having to learn the distribution of all the individual trajectories, we can instead just learn one velocity field representing their bulk evolution.. Our previous discussion [^46] Compare to Equation (49) in Section 3. A formal statement of how to combine flows is given in Appendix B.4. in Section 3 tells us how to do this: to determine the effective velocity $v_{t}^{*}(x_{t})$, we should take a weighted-average of all individual particle velocities $v^{[x_{1},x_{0}]}_{t}$, weighted by the probability that a particle at $x_{t}$ was generated by the pointwise flow $v^{[x_{1},x_{0}]}$. The final result is474747 An alternate way of viewing this result at a high level is: we start with pointwise flows $v^{[x_{1},x_{0}]}$ which transport delta distributions: $\delta_{x_{1}}\overset{v^{[x_{1},x_{0}]}}{\hookrightarrow}\delta_{x_{0}}.$ (62) And then Equation (64) gives us a fancy way of “averaging these flows over $x_{1}$ and $x_{0}$”, to get a flow $v^{*}$ transporting $q=\mathop{\mathbb{E}}_{x_{1}\sim q}[\delta_{x_{1}}]\overset{v^{*}}{\hookrightarrow}\mathop{\mathbb{E}}_{x_{0}\sim p}[\delta_{x_{0}}]=p.$ (63)
@@ 4.4 A Simple Choice of Pointwise Flow
We need an explicit choices of: pointwise flow, base distribution $q$, and coupling $\Pi_{q,p}$. There are many simple choices which would work484848 Diffusion provides one possible construction, as we will see later in Section 4.6..
@@
which simply linearly interpolates between $x_{1}$ and $x_{0}$ (and corresponds to the choice made in Liu et al. (2022a)). In Figure 5 we visualize a marginal flow composed of linear pointwise flows, the same annular base distribution $q$ of Figure 4, and target distribution equal to a point-mass ($p=\delta_{x_{0}}$) [^49] A marginal distribution with a point-mass target distribution – or equivalently the average of pointwise flows over the the base distribution only – is sometimes called a (one-sided) conditional flow (Lipman et al., 2023). .
@@ 4.5 Flow Matching
    Compute the point $x_{t}$ deterministically, by running515151 If we chose linear pointwise flows, for example, this would mean $x_{t}\leftarrow tx_{1}+(1-t)x_{0}$, via Equation (66). the pointwise flow $v^{[x_{1},x_{0}]}$ starting from point $x_{1}$ up to time $t$.
@@
To sample from the trained model (our estimate of $v_{t}^{*}$), we first sample a source point $x_{1}\sim q$, then transport it along the learnt flow to a target sample $x_{0}$. Pseudocode listings  4 and  5 give the explicit procedures for training and sampling from flow-based models (including the special case of linear flows for concreteness; matching Algorithm 1 in Liu et al. (2022a).).
@@ 4.6 DDIM as Flow Matching [Optional]
    which both generate the trajectory525252 See Appendix B.6 for details on why  (70) and (71) are equivalent along their trajectories.:
@@
This claim is straightforward to prove (see Appendix B.5), but the implication is somewhat surprising: we can recover the DDIM trajectories (which are not straight in general) as a combination of the straight pointwise trajectories in Equation (72). In fact, the DDIM trajectories are exactly equivalent to flow-matching trajectories for the above linear flows, with a different scaling of time ($\sqrt{t}$ vs. $t$)535353 DDIM at time $t$ corresponds to the linear flow at time $\sqrt{t}$; thus linear flows are “slower” than DDIM when $t$ is small. This may be beneficial for linear flows in practice (speculatively).. {mdframed}[nobreak]
@@ 4.7 Additional Remarks and References [Optional]
  • What we called pointwise flows are also called two-sided conditional flows in the literature, and was developed in Albergo and Vanden-Eijnden (2022); Pooladian et al. (2023); Liu et al. (2022a); Tong et al. (2023).
@@ 5 Diffusion in Practice
One advantage of the VAE interpretation is, it gives us an estimate of the data likelihood under our generative model, by using the standard Evidence-Based-Lower-Bound (ELBO) for VAEs. This allows us to train diffusion models directly using a maximum-likelihood objective. It turns out that the ELBO for the diffusion VAE reduces to exactly the L2 regression loss that we presented, but with a particular time-weighting that weights the regression loss differently at different time-steps $t$. For example, regression errors at large times $t$ (i.e. at high noise levels) may need to be weighted differently from errors at small times, in order for the overall loss to properly reflect a likelihood.585858 See also Equation (5) in Kadkhodaie et al. (2024) for a simple bound on KL divergence between the true distribution and generated distribution, in terms of regression excess risks. The best choice of time-weighting in practice, however, is still up for debate: the “principled” choice informed by the VAE interpretation does not always produce the best generated samples595959 For example, Ho et al. (2020) drops the time-weighting terms, and just uniformly weights all timesteps.. See Kingma and Gao (2023) for a good discussion of different weightings and their effect.
@@ B.1 KL Error in Gaussian Approximation of Reverse Process
$=F(0)+0.5\sigma^{2}(F^{\prime\prime}(0)-F^{\prime}(0)^{2})+O(\sigma^{4})$ (by Line 103)
@@ B.6 Two Pointwise Flows for DDIM give the same Trajectory
We want to show that pointwise flow 71:
@@ B.8 Proof Sketch of Claim 2
Now continuing from Line 129,